OpenStudy (anonymous):

Write the slope-intercept form of the equation of the line which passes through the points (1, 6) and (2, 9). A. y = 2x + 3 B. y = 3x + 3 C. y = 6x + 9 D. y = 3x + 6

OpenStudy (anonymous):

so the slope of the line is 3, so we've narrowed it down to B and D. then what you do is y - 6 = 3(x - 1)

OpenStudy (anonymous):

\[y = 3x - 3 + 6\] \[y = 3x + 3\] so obtion B

OpenStudy (anonymous):

Wow thanks a lot especially in explaining

OpenStudy (anonymous):

also, to figure out the slope, I did \[\frac{changeiny}{changein x}\] so here i did \(\frac{9-6}{2-1}\) = 3